OpenStudy (anonymous):

Equation of a circle So I'm giving an equation of a circle x^2 + y^2 = 1 but I only want 1/4 of the circle that's located on the first quadrant. My question is, what's the equation for the 1/4 piece of a circle that's located on the first quadrant.

OpenStudy (anonymous):

The top semicircle is found by solving for y: x^2 + y^2 = 1 y^2 = 1 - x^2 y = sqrt(1-x^(2) Note, that y - - sqrt(1-x^2) will be the lower half of the semicircle. Now that you have the top half of the semicicle, which is in quadrants I and II. To find that part of the seicircle that lies in quadrant I, just limit the domain to [0,1]. Becuase the entire top semicircle is [-1,1]. So yo are looking for y = sqrt(1-x^2) on [0,1].

OpenStudy (anonymous):

So there's no equation for 1/4 of a circle that's on the first quadrant?

OpenStudy (anonymous):

The circle equation will be the same it is only the limits that you need to specify in the first quarter the coordinates are both positive and range of values for x and y is from 0 to 1. so by specifying the domain and range in that quarter and putting the same equation you have specified what you want.